An aqueous solution containing 5.22 g of lead(ll) nitrate is added to an aqueous solution containing 5.19 g of potassium chloride to generate solid lead(ll) chloride and potassium nitrate. Write the balanced chemical equation for this reaction. Be sure to include all physical states. Tip: If you need to clear your work and neset the equation, click the button that looks like two arrows. What is the limiting reactant? O ad(ll) nitrate O potassium chloride The percent yield for the reaction is 90.3%, how many grams of precipitate were recovered? Number How many grams of the excess reactant remain? Number
Answer #1

Pb(NO3)2 + 2KCl -------------> PbCl2 + 2KNO3

331.2 g Pb(NO3)2 reacts with 2 x 74.55 g KCl

5.22 g Pb(NO3)2 reacts with 5.22 x 2 x 74.55 / 331.2 = 2.35 g KCl

but we have 5.19 g KCl so KCl is exess reagent and Pb(NO3)2 is limiting reagent.

limiting reagent = lead (II) nitrate

331.2 g Pb(NO3)2 forms 271.8 g PbCl2

5.22 g Pb(NO3)2 forms 5.22 x 271.8 / 331.2 = 4.28 g

% yield = 90.3

90.3 = (X / 4.28) x 100

X = 3.86 g

mass of precipitate = 3.86 g

mass of exess reagent = 5.19 - 2.35 = 2.84 g

exess reagent left = 2.84 g